OpenStudy (anonymous):

HElp ANYONE?? What transformation has changed the parent function f(x) = log3x to its new appearance shown in the graph below? logarithmic graph passing through point negative 1, 3 f(x + 4) + 2 f(x − 2) − 2 f(x + 2) + 2 f(x − 4) − 2

OpenStudy (anonymous):

guess you have to compute \[\log_3(x+4)+2\] at \(x=-1\) and see if you get \(3\)

OpenStudy (anonymous):

oh wow get it on the first try!! i figured it would be C, but it is not \[\log_3(-1+4)+2=\log_3(3)+2=1+2=3\] what a surprise!

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Can you help me with another question?

OpenStudy (anonymous):

So what exactly was the answer?